OpenStudy (anonymous):

sin^2(5) + sin^2(10) +............+sin^2(90)/(1+cos^2(5) + cos^2(10)......cos^2(90)) = ?

OpenStudy (anonymous):

Did Nt understood that

OpenStudy (anonymous):

Try to use the fact that cos( 90 - A) = sinA.

OpenStudy (anonymous):

OK i used it and got 1/2 but the answer should be 1

OpenStudy (anonymous):

There you would see that, waiiittt. The numerator would be equal to the denominator. You should be left with 1, I guess.

OpenStudy (ujjwal):

Add a sin^2 (0) to numerator. And in denominator, 1=cos^2 (0). Now, Sin^2 theta = cos^2 (90-theta) Proceed yourself further.

OpenStudy (ujjwal):

Convert all to cos

OpenStudy (anonymous):

should be add simply sin^2 (0)

OpenStudy (ujjwal):

What is the value of sin0 ?

OpenStudy (anonymous):

ok i get it thanzzzz @ujjwal

OpenStudy (anonymous):

Yeah. Look, sin^2 5 = cos^85. sin^ 10 = cos^80. Hence all terms equal.

OpenStudy (anonymous):

@ujjwal can u do chemistry

OpenStudy (ujjwal):

@shakir Don't tag me in your questions. And don't send me messages providing links to your questions.. you will get help but you must wait for a while.. And i am not that good in chemistry.

OpenStudy (anonymous):

ok